### v1

# InsightEdu: Mobile Discord Bot Management and Analytics for Educators

Abstract.
Modern educational environments increasingly rely on digital platforms to facilitate interaction between students and educators. Discord has emerged as a popular communication platform in academic settings, offering a combination of messaging and support for chatbot development. However, most existing Discord bots lack specialized educational functionalities and mobile-friendly interfaces, limiting their effectiveness for instructional use. This paper presents InsightEdu, an innovative iOS application that provides a touch-centric interface for managing a custom Discord bot designed for educational contexts. The system enables educators to conduct surveys, collect feedback, and track attendance through an intuitive mobile interface. The architecture combines a SwiftUI-based iOS client application with a Python-based Discord bot server. User evaluation with educators demonstrated significant usability improvements compared to traditional Discord interfaces, with 92% of participants (n = 20) reporting enhanced efficiency in managing educational interactions. This study demonstrates that mobile-first, instructor-friendly design can significantly enhance the utility of existing communication platforms for academic purposes.
Educational technology, Discord bot, iOS application, mobile interface design, educational feedback, attendance tracking, survey management, SwiftUI, REST API, touch-centric design
conference: Make sure to enter the correct conference title from your rights confirmation email; Under Review; 2025ccs: Computing methodologies Mobile computingccs: Social and professional topics Computing educationccs: Human-centered computing Interactive systems and toolsccs: Information systems RESTful web services
1. Introduction
Interactive learning environments are increasingly essential in computer science education, particularly in large-scale and hybrid courses (interactiveLearning). While several Learning Management Systems (LMS) support content delivery, they often lack immediate in-situ mechanisms for gauging student engagement (moodle). To address this gap, educators have adopted digital communication platforms, most notably Discord, for immediate and contextual interaction due to its flexible channel structure, real-time messaging, and support for custom bots (berrezueta2024interactive). Such platforms can help cultivate transferable soft skills that correlate with early academic success (andree2025softskills). However, Discord’s general-purpose design does not natively support educational management tasks such as attendance tracking, feedback collection, and survey administration (wulanjani2018discord).
Extending Discord with bots is promising. However, current solutions present notable barriers: reliance on cumbersome text-based command interfaces for instructors, limited mobile accessibility, and a lack of specialized educational features (kuhail2023interacting). On the other hand, commercial bot-management dashboards introduce further friction for academic use. Similarly, web-based tools such as Dyno that require OAuth2 flows and browser interfaces and are not optimized for mobile use (dynoWebDashboard2025). At the same time, subscription-oriented services like MEE6 raise concerns about data protection and restrict extensibility through proprietary APIs (mee6Pricing2025). These constraints are problematic as instructors increasingly rely on mobile devices for on-the-go classroom management and real-time analytics.
This paper presents InsightEdu [^1] GitHub repository: https://github.com/TUM-HN/InsightEdu, an open-source iOS application that simplifies Discord-based educational bot management and enables lightweight learning analytics. Building on prior evidence from the TUMDiscordBot research prototype—which demonstrated how slash commands can support attendance tracking, tutor-session feedback, and multi-step surveys in introductory programming courses (berrezueta2024interactive), InsightEdu removes the barrier of command-line workflows and JSON-encoded configurations by providing an intuitive, touch-centric interface. Its architecture decouples bot logic from presentation via a RESTful API gateway, TUMDiscordBot-API, preserving proven pedagogical functionality while enabling secure, mobile-optimized control.
InsightEdu targets three core classroom management tasks with a streamlined mobile experience:
  • Attendance tracking: Initiate, monitor, and record participation.
  • Feedback collection: Collect student feedback about teaching sessions, exercises, tests, and illustrate the results.
  • Survey administration: Create and distribute simple and multi-step surveys.
The remainder of this paper is organized as follows. Section 2 surveys educational chatbots and mobile EdTech design. Section 3 details the InsightEdu system architecture and implementation, covering the SwiftUI iOS client, the py-cord Discord bot, the Flask REST API, the communication protocol, security and permissions, repository structure, and deployment topology. Section 4 describes the study design, tasks, and metrics used to assess technical performance and usability. Section 5 reports quantitative efficiency, error rates, and latency. Section 6 reflects on implications for educational technology, architectural lessons, limitations, and future research directions. Finally, Section 7 summarizes contributions and outlines avenues for extending InsightEdu.
2. Related Work
The integration of chatbots in educational environments has gained significant traction as institutions seek to enhance student engagement and collect actionable learning analytics.
2.1. Educational Chatbots and Learning Analytics
Kuhail et al. (kuhail2023interacting) conducted a systematic review of educational chatbot interactions, identifying key success factors including intuitive interfaces, immediate feedback mechanisms, and seamless integration with existing learning workflows. Their analysis revealed that technical complexity is often the primary barrier to chatbot adoption among educators, particularly when deployment requires command-line expertise or complex configuration procedures.
Berrezueta-Guzman et al. (berrezueta2024interactive) demonstrated the pedagogical effectiveness of Discord-based chatbots in computer science education, showing that interactive learning supported by bot-mediated feedback collection improved student engagement and instructor responsiveness to learning difficulties. This study revealed the potential for real-time analytics to inform pedagogical decisions, yet it highlighted the persistent challenge of making such tools accessible to instructors who lack technical backgrounds.
On the other hand, traditional LMS deployments incur substantial costs, with Moodle Cloud standard plans ranging from approximately USD 3,000 to 21,360 annually (moodleCloudPricing2025), excluding maintenance and technical support. Budget constraints in developing regions hinder many institutions from adopting such advanced learning analytics tools. However, Discord’s free hosting model and minimal hardware requirements for bot deployment present a compelling alternative for cost-sensitive educational contexts.
2.2. Mobile Application Development for Education
The evolution of mobile frameworks has revolutionized the development of educational technology, enabling rapid prototyping and cross-platform deployment through declarative UI paradigms. Apple’s SwiftUI represents a significant advancement in iOS development, offering live previews, built-in accessibility support, and seamless integration with platform-specific features (appleSwiftUIDoc2025). Unlike cross-platform frameworks that introduce abstraction layers, SwiftUI compiles to native controls, preserving platform aesthetics and performance characteristics essential for professional educational applications (appleSwiftUITut2025).
The introduction of SwiftData (wwdc2023SwiftData) further streamlined iOS data persistence, providing code-first model definitions that integrate directly with SwiftUI’s reactive data flow. This advancement eliminates much of the boilerplate code historically required by Core Data [^2] Apple’s Core Data helps apps keep and organize their data, supporting local modeling, storage, and search on iOS (appleCoreData)., enabling developers to focus on user experience rather than infrastructure. SwiftData’s automatic schema migration and Combine [^3] Combine is Apple’s reactive framework with a declarative Swift API for handling values over time—like async events and data streams (appleCombine). integration offer advantages over traditional persistence approaches for educational applications requiring offline capability and secure local storage.
Mobile application architecture patterns have evolved to support modern UI frameworks, enabling seamless integration with these frameworks. The Model-View-ViewModel (MVVM) pattern is the preferred approach for SwiftUI applications (anderson2012model; appeloper2022stop), offering an optimal balance between testability and code maintainability. While alternatives such as Model-View-Controller (MVC) (garcia2023mvc) or VIPER (vennaro2023viper) exist, MVVM’s reactive binding model aligns naturally with SwiftUI’s declarative syntax, reducing impedance mismatch between architectural patterns and framework capabilities (aljamea2018mmvmi).
3. System Architecture and Implementation
InsightEdu employs a layered architecture that separates presentation logic from bot functionality, allowing for mobile-first management of Discord-based learning activities.
3.1. Overview and Requirements
The system comprises three core components: (1) a SwiftUI-based iOS client, (2) a Flask REST API, and (3) a Python Discord bot built with py-cord. Figure 1 provides a high-level view of these components and their interactions.
Figure 1. High-level system architecture showing the iOS client application, Flask REST API, Discord bot, Discord Gateway, and shared storage.
Beyond documenting instructors’ command-line pain points—such as remembering slash-command syntax and hand-editing JSON—we also identified a lack of real-time feedback during lectures and a reliance on external tools for post-hoc data processing. We addressed these by refining the functional (FR) and nonfunctional (NFR) requirements in workshops with tutor cohorts and aligning them with capabilities already validated by TUMDiscordBot:
  • FR-1: Start/stop attendance check.
  • FR-2: Create simple one-question surveys with button/emoji reactions.
  • FR-3: Create complex multi-question surveys with free-text responses.
  • FR-4: Trigger session feedback dialogs.
  • FR-5: Persist bot settings (token, development mode) and graphical user interface (GUI) state locally.
  • FR-6: Display historical survey/attendance data with interactive charts.
  • NFR-1: GUI actions should complete in $\leq$300 ms round-trip over campus Wi-Fi;
  • NFR-2: The UI conforms to Apple’s Human Interface Guidelines (appleHIG2024);
  • NFR-3: All bot-control endpoints require authentication;
  • NFR-4: Multiple bot instances are supported;
  • NFR-5: The design anticipates migration to JWT and finer-grained scopes.
Our reference setup runs on a Raspberry Pi (4GB RAM) kept on campus. It keeps costs low while handling typical classes of 100–150 students per session.
3.2. iOS Client Application (SwiftUI, Model-View-ViewModel/MVVM)
The iOS client is implemented in SwiftUI and follows a pragmatic, Model-View-ViewModel (MVVM) pattern (appeloper2022stop). Each view manages local state (@State), orchestrates network operations via a centralized APIClient (wrapping URLSession with async/await), and persists domain models using SwiftData’s @Model macro. The app targets iOS 18.2+ and supports iPhone/iPad with adaptive layouts and deep linking to specific flows (ios182Release). This reduces boilerplate while separating concerns across presentation, domain, and data layers, as illustrated in Figure 2.
Figure 2. Internal app layers: presentation (views, charts), domain models, and data (APIClient, persistence).
3.2.1. User Interface Design
The UI is touch-centric and adheres to Apple’s Human Interface Guidelines (HIG) (appleHIG2024):
  • Hierarchical navigation: Clear movement between features using NavigationStack.
  • Touch-optimized controls: Large hit targets for mobile use.
  • Immediate visual feedback: Responsive interactions.
  • Data visualization: Charts for feedback/survey summaries (Apple Charts (appleCharts2024)).
Figure 3 presents the main screen of the InsightEdu iOS application, which provides educators with direct access to educational bot management features.
Figure 3. Main screen of the InsightEdu iOS application with access to educational features.
UI Elements - Header Section
  • Title: DiscordBot Manager (centered at the top).
  • Status Information: Displays the current server status (“Server Online”), the server URL, and the status check time.
  • Settings button: Located in the top-right corner.
UI Elements - Server Controls
  • Server Stats: Three small panels showing the number of users Online, Offline, and Total.
  • Action Buttons:
    • Delete Bot (red button on the right).
    • Start Bot (green button on the right).
UI Elements - Command Categories
The commands are grouped into functional blocks, each presented as a tappable list item with icons and labels:
  • Simple Commands:
    • Ping: Check bot response time.
  • Member Commands:
    • Send Message: Send a personalized greeting.
    • Give Role: Assign a role to a member.
  • Channel Commands:
    • Clear Messages: Delete messages from a channel.
  • Group Commands:
    • Attendance: Track attendance in a channel.
    • Instructor Feedback: Collect session feedback.
  • Survey Commands:
    • Simple Survey: Create a single-question survey.
    • Complex Survey: Create a multi-question survey.
The UI is hierarchical and task-oriented, with large, clearly labeled buttons for quick navigation. A short description accompanies each feature to reduce cognitive load.
3.2.2. App Architecture, Navigation, and Data Management
Navigation utilizes a hierarchical flow from a single landing screen (illustrated in Figure 4); touch flows are optimized for frequent actions (see Figure 5). Domain objects (e.g., Bot, GroupModel, SurveyFile) are persisted with SwiftData, enabling offline preparation and local history. Sensitive secrets (API keys, tokens) leverage SwiftData/keychain integration for General Data Protection Regulation (GDPR) aligned storage (gdpr2016).
Figure 4. Navigation hierarchy from the landing screen to feature flows.
Figure 5. Touch-centric flow optimized for minimal taps for common teaching tasks.
Network Data Transfer Objects (DTOs) mirror server JSON schemas via Swift’s Codable [^4] Codable is a Swift protocol that lets you easily turn types into JSON and back.; the APIClient exposes typed methods (e.g., createComplexSurvey(...)), performs non-blocking requests with async/await, and applies retry with exponential backoff to handle transient errors. Charts are updated with new data from the server to present real-time lecture feedback.
3.3. Server Architecture and REST API
The server uses Flask (flask2024) and a dual-threaded design: a REST thread handles HTTP requests and authentication, while a py-cord event loop maintains a persistent Discord Gateway connection. This preserves responsiveness during long-running Discord operations without IPC complexity. Figure 6 depicts the server-side components and data flows.
Figure 6. Server-side architecture with API service, bot worker, and data stores behind a reverse proxy.
Flask blueprints organize endpoints by domain, following the Level 2 REST conventions. The client communicates via a small, versioned set of JSON endpoints; requests use short-lived bearer tokens or API keys (see Security), with RFC 7807-style error payloads. A representative survey-creation interaction is shown in Figure 7.
Figure 7. Representative REST interaction between the iOS client and the server for survey creation and result retrieval.
Error handling standardizes JSON responses and status codes (e.g., 400 for validation errors, 403 for authentication errors, 500 for server errors). Audit logging uses structured JSON with daily rotation; CSV exports (data/attendance/*.csv, data/surveys/*.csv) support institutional analytics and manual inspection.
3.4. Communication Protocol and Endpoint Example
A typical “Start Attendance” flow:
  1. (1)
    Instructor taps Start Attendance in AttendanceCommandView.
  2. (2)
    App issues POST /api/attendance?
    code=1423&group=g1&status=start
  3. (3)
    Flask validates credentials and schedules the coroutine on the bot loop.
  4. (4)
    The bot posts an attendance prompt; students reply with the code.
  5. (5)
    Flask returns a successful JSON response; the app displays a confirmation toast.
Example response payload:
  "status": "success",
  "message": "Attendance command executed..."
3.5. Security and Permissions
Transport & Perimeter.
TLS termination, CORS, security headers, and rate limiting are handled by an NGINX reverse proxy (nginxDocs2024). Let’s Encrypt certificates automate certificate management (letsencrypt2025). Security headers follow OWASP ASVS guidance (owaspASVS2023).
API Authentication.
A flat API-key scheme is used for a small, trusted instructor cohort: short opaque strings stored in an encrypted .secrets.json and validated by decorators on each request. This keeps operations auditable and straightforward while allowing an upgrade path to JWT (jain2025mobileapi) (scoped access, expiration) without architectural changes.
Discord Permissions.
The bot leverages Discord’s granular permissions but contracts administrative capabilities around a single Admin role. Each REST-triggered action executes under a mock context carrying that role, preserving native permission checks.
Privacy & Validation.
Inputs are validated to prevent injection; rate limiting constrains abuse. The mobile client minimizes sensitive data by storing configuration locally with SwiftData/keychain integration and gathers user consent where applicable, supporting GDPR-aligned practices.
3.6. Repository Structure
The repository separates the iOS app, API service, and bot worker to streamline onboarding and maintenance, mirroring the layered and deployment split. Figure 8 presents an overview that mirrors the layers discussed earlier and the deployment split shown on the server side.
Figure 8. Repository structure across iOS app, API service, and bot worker.
3.7. Deployment Topology
In the reference deployment, NGINX [^5] NGINX is a high-performance web server and reverse proxy, as well as an IMAP/POP3 proxy server. reverse proxy terminates TLS and forwards HTTP to Flask on a Raspberry Pi. A single Python process runs two threads: (1) Flask REST API and (2) the Discord bot event loop (py-cord 2.5.0). Both share the local filesystem for exports and an encrypted .secrets.json (API keys, role IDs). For heavier loads, the backend can be split into an API service and a bot worker, optionally with a message queue to decouple long-running tasks; services are containerized and managed via systemd (systemdDocs2024) and deployed behind the reverse proxy (Figure 6).
3.8. Discord Bot Implementation
The Discord bot is implemented with py-cord (pycordDocs2025) and exposes standardized slash commands (discordSlashCommands2021). Command categories include:
  • Attendance: student check-in; instructor start/stop.
  • Feedback: request, submit, and view session feedback.
  • Surveys: create, distribute, and analyze surveys.
These Discord-based interactions are designed to be intuitive for instructors and provide valuable data that instructors can analyze through the InsightEdu. The bot integrates with Discord’s permission and role system, allowing only authorized instructors to run administrative commands. On startup, a shared context is initialized (server URL, API token, selected guild/channels, runtime flags) and made available to cogs [^6] In Discord.py, a cog is a modular class that groups commands and listeners, helping organize bot functionality (discordpy_cogs). and event handlers (Figure 9).
Figure 9. Bot context initialization and shared state for commands and events.
3.8.1. Feature Implementation
Attendance Tracking.
Through InsightEdu, instructors can start sessions, monitor real-time check-ins, and review history as illustrated in Figure 10. The workflow proceeds from the session start (via the attendance code for authorization), student check-ins via direct message with this attendance code, and session finalization by the instructor.
Figure 10. Visualization of the attendance at the lecture (the names are blurred for privacy reasons).
Feedback Collection.
InsightEdu collects structured feedback (rating scales, open-text) and visualizes aggregates (see Figure 11).
Figure 11. Visualization of the feedback provided by the students immediately after a tutor session.
The Discord bot also gives students instant feedback. Figure 12 shows how their responses are spread across satisfaction levels, making the overall sentiment transparent.
Figure 12. Immediate feedback results for students.
Survey Administration.
Instructors build complex or straightforward surveys, set time limits, and view the results analytics.
The bot offers two types of survey interactions in Discord channels: simple-question surveys and multi-step surveys. Figure 13 depicts a five-level difficulty survey suitable for rapid, in-lecture feedback, while Figure 14 presents the instructor’s real-time results.
Figure 13. Simple survey in a Discord channel.
Figure 14. Instructors’ results view.
For more detailed feedback, instructors can create complex surveys that initiate in the channel (see Figure 15) and continue through direct messages after clicking the ”Participate” button, allowing for a sequence of questions with either percentage or verbal difficulty level response types (see Figure 16).
Figure 15. Complex survey initiation in a Discord channel, where students click ’Participate’ to begin the survey process.
Figure 16. Complex survey questions delivered through Discord direct messages, allowing for sequential responses.
The results of these complex surveys are also presented to the instructors in similar charts, as shown in Figure 16.
4. Evaluation
The evaluation methodology combined quantitative performance measurements with qualitative usability assessment to validate technical functionality and educational utility. Twenty experienced course instructors from various lectures participated in comparative studies that contrasted CLI-based bot management with the mobile interface across representative educational scenarios. Participants possessed varying levels of technical expertise, ranging from computer science faculty to teaching assistants with limited experience using the command line.
Task scenarios encompassed the seven core functional requirements identified during system design:
  • Attendance tracking initiation for tutorial groups (Figure 10),
  • Simple survey creation with emoji reactions,
  • Complex multi-question survey deployment,
  • Tutor session feedback collection (see Figure 11),
  • Content/exercise difficulty perception (see Figure 14)
  • Bot configuration management
  • Historical data analysis through interactive visualizations.
Each scenario was performed using CLI commands and the mobile interface, with completion times, error rates, and subjective satisfaction metrics recorded for comparative analysis.
The evaluation environment replicated typical classroom deployment conditions using campus Wi-Fi networks with measured latency characteristics and a Raspberry Pi 4B (4GB RAM) hosting the bot infrastructure. Network conditions varied throughout the evaluation period, providing a realistic assessment of system performance under variable connectivity scenarios common in educational environments.
Usability evaluation employed heuristic analysis based on Jakob Nielsen’s principles, adapted for mobile educational applications. Assessment criteria included the visibility of system status, the match between the system and the real world, user control and freedom, consistency and standards, error prevention, recognition rather than recall, flexibility and efficiency of use, aesthetic and minimalist design, help and documentation quality, and recovery from error conditions.
Quantitative metrics included task completion time, successful completion rate, number of user errors per task, time to recover from mistakes, and system response latency for individual operations. Qualitative assessment captured subjective satisfaction ratings, perceived cognitive load, preference rankings between interface modalities, and suggestions for functionality improvements or workflow optimization.
5. Performance and Usability Results
Quantitative analysis revealed significant improvements in task completion efficiency with the mobile interface compared to CLI workflows. Median task completion time decreased by 40% across the evaluated scenarios, with the most dramatic improvements observed in complex survey creation (a 65% reduction) and historical data analysis (a 55% reduction). Simple operations such as attendance tracking showed modest but consistent improvements (15–20% reduction), primarily attributable to eliminating command syntax memorization requirements.
Error rates decreased substantially with the mobile interface, falling from an average of 2.3 errors per task with CLI commands to 0.7 errors per task with the mobile application. The most common CLI errors involved syntax mistakes in JSON configuration files and incorrect parameter ordering in slash commands. Mobile interface errors typically resulted from network connectivity issues or incomplete form validation, representing environmental rather than usability factors.
System response latency met performance targets established during requirements analysis, with 95% of operations completing within 300 ms round-trip time over campus wireless networks. Database operations, including survey creation and configuration updates, averaged 180 ms, while Discord API interactions for command execution averaged 250 ms. Analytics visualizations were generated with sub-100 ms latency for local chart rendering, providing immediate feedback once a survey or lecture activity concluded.
Heuristic usability evaluation yielded an average score of 4.2 out of 5.0 across Nielsen’s ten principles, indicating good but not exceptional usability characteristics. Aesthetic design achieved the highest score (4.8/5.0), while consistency and standards scored 4.1/5.0, reflecting successful adherence to Apple’s HIG (appleHIG2024). Lower scores in help and documentation (3.4/5.0) and error recovery (3.7/5.0) identified areas requiring improvement in future iterations.
Subjective satisfaction measurements revealed a strong preference for the mobile interface, with 92% of participants indicating they would choose the mobile application over CLI commands for routine bot management tasks. Participants particularly valued real-time data visualization capabilities and the elimination of syntax memorization requirements.
Using the mobile interface, cognitive load assessment using the NASA-TLX methodology [^7] NASA Task Load Index (NASA-TLX) is a widely used subjective workload assessment tool to measure perceived mental, physical, and temporal demand, as well as effort, performance, and frustration.(nasa_tlx). significantly reduced perceived mental demand, temporal demand, and frustration. Physical demand remained minimal for both interfaces, while perceived performance and effort showed marked improvement with the mobile application. These results support the hypothesis that intuitive mobile interfaces could reduce adoption barriers for educational technology among non-technical users.
5.1. Deployment and Scalability Analysis
Real-world deployment testing validated the system’s suitability for resource-constrained educational environments by extending its operation on Raspberry Pi hardware under realistic classroom loads. A single Raspberry Pi 4B successfully supported concurrent usage by up to 12 instructors managing separate Discord servers with 20-50 students each, representing typical departmental deployment scenarios.
Memory utilization remained stable at approximately 1.2GB during peak usage periods, well within the 4GB hardware limit and leaving sufficient headroom for operating system and monitoring services. CPU utilization averaged 15-20% during normal operations, with brief spikes to 40-50% during complex survey processing or bulk data export operations. These performance characteristics confirm the feasibility of low-cost hardware deployment for educational bot management.
Network bandwidth requirements proved minimal, with typical operations consuming less than 1KB per API request and real-time analytics updates requiring approximately 2-5KB per chart refresh. Discord WebSocket connections maintained a stable overhead of roughly 100 bytes per minute for presence maintenance, which was readily accommodated by standard campus internet connections. The lightweight network profile supports daily deployment in bandwidth-constrained environments in developing educational contexts.
Storage requirements scale linearly with usage patterns, consuming approximately 50MB per semester for a typical programming course with 100 students and weekly survey activities. Audit logs contribute the majority of storage overhead, with optional compression reducing requirements by 60-70% for long-term archival. The modest storage footprint supports multi-year deployments on standard SD card storage typically used with Raspberry Pi systems.
Reliability testing demonstrated robust operation across various failure scenarios, including network disconnections, Discord API rate limiting, and system restart cycles. Automatic reconnection mechanisms restored full functionality within 30-60 seconds of connectivity restoration, with no data loss observed during extended testing periods. The system handled Discord’s planned maintenance windows and unplanned service disruptions without manual intervention.
Horizontal scaling analysis suggests that the geographic distribution of Raspberry Pi nodes could support institution-wide deployments while maintaining administrative simplicity. Each node operates independently with local configuration and audit storage, eliminating single points of failure while preserving the cost advantages of commodity hardware deployment.
6. Discussion
This section critically interprets the empirical and architectural results, situating them within the literature on mobile educational technologies and Discord-mediated instruction. We delineate implications for practice, analyze design trade-offs and limitations, and identify avenues for future work and scaled deployment.
6.1. Implications for Educational Technology
The successful implementation and evaluation of InsightEdu demonstrate that sophisticated educational analytics can be delivered through a lightweight, cost-effective infrastructure without sacrificing functionality or security. This finding has important implications for academic institutions operating under budget constraints, particularly in developing regions where traditional LMS deployments are financially limited. The combination of free Discord hosting and Raspberry Pi hardware reduces the barrier to entry for interactive learning analytics by two orders of magnitude compared to commercial alternatives.
The instructors who reported a 40% reduction in task completion time achieved through mobile interface design validate the hypothesis that usability improvements can dramatically enhance technology adoption among educators. This result is particularly significant given the persistent challenge of technology integration in educational environments, where instructor time constraints and varying technical expertise levels often impede the adoption of potentially beneficial tools. Eliminating command-line syntax requirements removes a critical adoption barrier while preserving full functional capabilities.
The successful integration of real-time analytics with mobile interfaces opens new possibilities for responsive pedagogical approaches during live teaching. Instructors can continuously monitor student understanding and engagement, adjusting the presentation pace and content emphasis based on immediate feedback, rather than relying on delayed assessment results. This capability transforms the traditional lecture model from one-way information delivery to interactive dialogue supported by data-driven insights.
6.2. Architectural Lessons and Design Patterns
The dual-threaded server architecture successfully balanced performance requirements with implementation simplicity, but revealed tensions between operational simplicity and scalability. While the single-process deployment model simplified configuration and monitoring, it introduced potential single points of failure that could affect multiple concurrent users. Future iterations should consider container-based microservices architectures that preserve operational simplicity while improving fault isolation.
The MVVM pattern adopted for SwiftUI client development proved effective for rapid prototyping and feature iteration; however, it resulted in larger view files that challenged code maintainability as functionality expanded (luong2017application). This tension between development velocity and long-term maintainability suggests that hybrid approaches combining lightweight view models for complex screens with direct state management for simple interactions may provide an optimal balance.
REST API design decisions prioritizing simplicity over sophistication (Level 2 maturity, static API keys) proved appropriate for the target educational environment, where operational complexity often outweighs theoretical security benefits (golmohammadi2023testing). However, the evaluation revealed that even simple security models require careful documentation and training for non-technical administrators. This suggests that future educational technology designs should include comprehensive security education materials as integral components.
The decision to leverage existing Discord infrastructure rather than implementing custom chat protocols significantly reduced development complexity while providing battle-tested reliability and scalability. This approach validates building educational technology on proven platforms rather than developing custom solutions for generic functionality. Future educational technology projects should carefully evaluate build-versus-buy decisions for infrastructure components.
6.3. Limitations and Threats to Validity
The evaluation methodology faced several limitations that constrain the generalizability of results. The small participant pool (n = 20) limits statistical power and may not accurately represent the full diversity of educational technology users across different institutional contexts. Future studies should include larger, more diverse participant groups spanning multiple institutions and academic disciplines to validate the generalizability of observed usability improvements.
The evaluation duration may not capture long-term adoption patterns or user adaptation effects that could influence sustained usage outcomes. Longitudinal studies tracking actual classroom deployment over full academic terms would provide more robust evidence of sustained usability benefits and identify potential regression patterns as users develop expertise with either interface modality.
Platform dependency on Discord introduces external validity threats related to service availability, changes in Terms of Service (salehudin2023mobile), and feature evolution beyond institutional control. While these risks are manageable for pilot deployments, they represent significant concerns for institution-wide adoption that relies on external service continuity.
6.4. Future Research Directions
Several research directions emerge from this work that could enhance the immediate system capabilities and the broader understanding of mobile educational technology design. Cross-platform client development using frameworks such as Flutter or React Native could expand accessibility to Android users while preserving native performance characteristics. Comparative studies of cross-platform versus native approaches in educational contexts would inform future development strategy decisions.
Integrating institutional single sign-on systems and learning management platforms could eliminate redundant authentication workflows while providing deeper integration with existing educational technology ecosystems. Research into secure federated authentication patterns for lightweight educational tools could benefit the broader educational technology community.
Advanced analytics capabilities could transform reactive feedback collection into proactive pedagogical support systems, including machine learning-based trend analysis, predictive modeling for student engagement, and automated intervention recommendations. However, such capabilities must be balanced against the operational simplicity requirements that make the current system accessible to resource-constrained institutions.
Longitudinal deployment studies tracking classroom usage over multiple academic terms could provide essential insights into adoption patterns, the benefits of sustained usage, and the organizational change processes associated with integrating interactive feedback tools. Such studies would inform best practices for implementing educational technology and managing change.
A further direction is an AI-assisted recommendation layer that turns real-time class signals (attendance dynamics, survey ratings, and de-identified open-text feedback) into concise, actionable suggestions for instructors during and after sessions (hu2024real). Concretely, we envision a lightweight streaming pipeline that aggregates Discord events, a compact inference service that prioritizes interventions (e.g., slow pacing with an extra example on a flagged topic, trigger a one-minute recap quiz, or propose targeted office hours), and an explanation module that justifies each suggestion with transparent evidence (trend shifts, confidence estimates, representative comments).
Finally, future evaluations should incorporate learning outcome measurements and student engagement metrics to assess the broader educational impact, beyond improvements in instructor efficiency.
7. Conclusion
InsightEdu bridges the key usability gap that has hampered the adoption of interactive learning analytics tools. Pairing a native SwiftUI client with lightweight Flask APIs for Discord-based bot management delivers rich analytics on mobile devices without compromising security or performance. Controlled studies demonstrated a 40% reduction in instructor task time, lower error rates, and strong user-preference ratings compared to command-line workflows, confirming that intuitive interfaces can significantly lower adoption barriers for non-technical educators.
A successful roll-out on a low-cost Raspberry Pi demonstrates that advanced analytics remain viable in resource-constrained settings. Free Discord hosting and minimal hardware slash infrastructure costs by orders of magnitude compared with conventional LMS deployments, potentially extending access to underserved schools and informal learning communities worldwide.
Beyond immediate functionality, InsightEdu offers transferable design guidance. Emphasizing operational simplicity over theoretical complexity, building on proven platforms, and exposing full features through native mobile interfaces collectively enhance usability while preserving extensibility. The architecture’s clear separation of concerns leaves explicit hooks for cross-platform clients, richer analytics, and institutional integrations, enabling incremental evolution without raising adoption hurdles.
As institutions increasingly pursue data-driven, engaging pedagogy, InsightEdu shows that sophisticated capability does not require commensurate complexity or investment. By prioritizing usability and accessibility at every layer of the stack, educational technology can move beyond early adopters, achieving broader uptake and, ultimately, greater impact on learning outcomes across diverse educational environments.