### v2

# InsightEdu: Mobile Discord Bot Management and Analytics for Educators

Abstract.
Modern educational environments increasingly rely on digital platforms to facilitate interaction between students and educators. Discord has emerged as a popular communication platform in academic settings, offering a combination of messaging and support for chatbot development. However, most existing Discord bots lack specialized educational functionalities and mobile-friendly interfaces, limiting their effectiveness for instructional use. This paper presents InsightEdu, a mobile-first interaction design and reference architecture for instructor-operated Discord bots. The system enables educators to conduct surveys, collect feedback, and track attendance through an intuitive mobile interface. The architecture combines a SwiftUI-based iOS client application with a Python-based Discord bot server. User evaluations with educators demonstrated significant usability improvements over traditional Discord interfaces, with 92% of participants (n = 20) reporting enhanced efficiency in managing educational interactions. This study demonstrates that mobile-first, instructor-friendly design can significantly enhance the utility of existing communication platforms for academic purposes.
Educational technology, Discord bot, iOS application, mobile interface design, educational feedback, attendance tracking, survey management, SwiftUI, REST API, touch-centric design
conference: Make sure to enter the correct conference title from your rights confirmation email; Under Review; 2025ccs: Computing methodologies Mobile computingccs: Social and professional topics Computing educationccs: Human-centered computing Interactive systems and toolsccs: Information systems RESTful web services
1. Introduction
Interactive learning environments are increasingly essential in computer science education, particularly in large-scale and hybrid courses (Krusche and Berrezueta-Guzman, 2023). While several Learning Management Systems (LMS) support content delivery, they often lack immediate in-situ mechanisms for gauging student engagement (Gamage et al., 2022). To address this gap, educators have adopted digital communication platforms, most notably Discord, for immediate, context-specific interaction due to its flexible channel structure, real-time messaging, and support for custom bots (Berrezueta-Guzman et al., 2024). Such platforms can help cultivate transferable soft skills that correlate with early academic success (Andree et al., 2025). However, Discord’s general-purpose design does not natively support educational management tasks such as attendance tracking, feedback collection, and survey administration (Wulanjani, 2018).
Extending Discord with bots is promising. However, current solutions present notable barriers: reliance on cumbersome text-based command interfaces for instructors, limited mobile accessibility, and a lack of specialized educational features (Kuhail et al., 2023). On the other hand, commercial bot-management dashboards introduce further friction for academic use. Similarly, web-based tools such as Dyno that require OAuth2 flows and browser interfaces and are not optimized for mobile use (Aquino et al., 2025). At the same time, subscription-oriented services like MEE6 raise concerns about data protection and restrict extensibility through proprietary APIs (MEE6 Community, ). These constraints are problematic as instructors increasingly rely on mobile devices for on-the-go classroom management and real-time analytics.
This paper presents InsightEdu [^1] GitHub repository: https://github.com/TUM-HN/InsightEdu, an open-source iOS application that simplifies Discord-based educational bot management and enables lightweight learning analytics. Building on prior evidence from the TUMDiscordBot research prototype—which demonstrated how slash commands can support attendance tracking, tutor-session feedback, and multi-step surveys in introductory programming courses (Berrezueta-Guzman et al., 2024), InsightEdu removes the barrier of command-line workflows and JSON-encoded configurations by providing an intuitive and touch-centric interface. Its architecture decouples bot logic from presentation via a RESTful API gateway, TUMDiscordBot-API, preserving proven pedagogical functionality while enabling secure, mobile-optimized control.
InsightEdu targets three core classroom management tasks with a streamlined mobile experience:
  • Attendance tracking: Initiate, monitor, and record participation.
  • Feedback collection: Collect student feedback about teaching sessions, exercises, tests, and illustrate the results.
  • Survey administration: Create and distribute simple and complex surveys.
The remainder of this paper is organized as follows. Section 2 surveys educational chatbots and mobile EdTech design. Section 3 details the InsightEdu system architecture and implementation. Section 4 describes the study design, tasks, and metrics used to assess technical performance and usability. Section 5 reports quantitative efficiency, error rates, and latency. Section 6 reflects on implications for educational technology, architectural lessons, limitations, and future research directions. Finally, Section 7 concludes the paper.
2. Related Work
The integration of educational chatbots has gained traction as a tool to enhance student engagement and collect learning analytics. Kuhail et al. (Kuhail et al., 2023) identified that technical complexity remains the primary barrier to chatbot adoption among educators. While Discord-based bots have proven effective in improving student-instructor interaction in computer science courses, existing solutions often require command-line expertise for deployment and configuration (Berrezueta-Guzman et al., 2024). Furthermore, traditional Learning Management Systems (LMS) like Moodle Cloud incur substantial annual costs (up to USD 21,360) (MoodleCloud, 2025). Discord’s free hosting model provides a cost-effective alternative for budget-constrained institutions.
Modern declarative UI frameworks like SwiftUI have revolutionized educational technology by enabling rapid prototyping with native performance and built-in accessibility (Apple Inc., 2025b, c). SwiftData further streamlines this by providing code-first data persistence that integrates with SwiftUI’s reactive data flow, replacing the high boilerplate requirements of legacy frameworks like Core Data (Apple Inc., 2023, 2025a).
Model-View-ViewModel (MVVM) is the preferred architectural pattern for SwiftUI applications (Magics-Verkman et al., 2023). Unlike MVC or VIPER, MVVM’s reactive binding model aligns naturally with declarative syntax, reducing the impedance mismatch between logic and presentation layers (Anderson, 2012; Aljamea and Alkandari, 2018).
3. System Architecture and Implementation
InsightEdu employs a layered architecture that separates presentation logic from bot functionality, allowing for mobile-first management of Discord-based learning activities.
3.1. Overview and Requirements
The system comprises three core components: (1) a SwiftUI-based iOS client, (2) a Flask REST API, and (3) a Python Discord bot built with py-cord.
Beyond documenting instructors’ command-line pain points, such as remembering slash, command syntax, and hand-editing JSON, we also identified a lack of real-time feedback during lectures and a reliance on external tools for post-hoc data processing. We addressed these by refining the functional (FR) and nonfunctional (NFR) requirements in workshops with tutor cohorts and aligning them with capabilities already validated by TUMDiscordBot:
  • FR-1: Start/stop attendance check.
  • FR-2: Simple surveys with button/emoji reactions.
  • FR-3: Complex multi-question surveys with free-text responses.
  • FR-4: Trigger session feedback dialogs.
  • FR-5: Persist bot settings (token, development mode) and graphical user interface (GUI) state locally.
  • FR-6: Display historical survey/attendance data with interactive charts.
  • NFR-1: GUI actions should complete in $\leq$300 ms round-trip over campus Wi-Fi;
  • NFR-2: GUI follows Apple’s Human Interface Guidelines;
  • NFR-3: All bot-control endpoints require authentication;
  • NFR-4: Multiple bot instances are supported;
  • NFR-5: The design anticipates migration to JWT and finer-grained scopes.
Our bot setup runs on a Raspberry Pi (4GB RAM) kept on campus. It keeps costs low while handling typical classes of 100–150 students per session.
3.2. User Interface Design
The UI is touch-centric and adheres to Apple’s Human Interface Guidelines (HIG) (Apple Inc., 2024a):
  • Hierarchical navigation: Clear movement between features using NavigationStack.
  • Touch-optimized controls: Large hit targets for mobile use.
  • Immediate visual feedback: Responsive interactions.
  • Data visualization: Charts for feedback/survey summaries (Apple Charts (Apple Inc., 2024b)).
Figure 1 presents the main screen of the InsightEdu iOS application, which provides educators with direct access to educational Discord bot management features. Among its design we have:
UI Elements - Header Section
  • Title, Status Information, and Settings
UI Elements - Server Controls
  • Server Stats, Action Buttons (Delete Bot, Start Bot).
UI Elements - Command Categories The commands are grouped into functional blocks:
  • Simple Commands: Ping:
  • Member Commands: Send Message and Give Role
  • Channel Commands: Clear Messages
  • Group Commands: Attendance and Instructor Feedback
  • Survey Commands: Simple Survey and Complex Survey
Figure 1. Main screen of the InsightEdu iOS application.
The UI is hierarchical and task-oriented, with large, clearly labeled buttons for quick navigation. Additionally, a short description accompanies each feature to reduce cognitive load.
3.3. Server Architecture and REST API
The server uses Flask (Pallets Project, 2024) and a dual-threaded design, a REST thread handles HTTP requests and authentication, while a py-cord event loop maintains a persistent Discord Gateway connection.
Flask blueprints organize endpoints by domain, following the Level 2 REST conventions. The client communicates via a small, versioned set of JSON endpoints; requests use short-lived bearer tokens or API keys, with RFC 7807-style error payloads.
Error handling standardizes JSON responses and status codes. Audit logging uses structured JSON with daily rotation; CSV exports (data/attendance/*.csv, data/surveys/*.csv) support institutional analytics and manual inspection.
3.4. Discord Bot Implementation
The Discord bot is implemented with py-cord (Pycord Development Team, 2025) and exposes standardized slash commands (Discord Inc., 2021). Command categories include:
Attendance Tracking.
Through InsightEdu, instructors can start sessions, monitor real-time check-ins, and review history. The workflow proceeds from the session start (via the attendance code for authorization), student check-ins via direct message with this attendance code, and session finalization by the instructor.
Feedback Collection.
InsightEdu collects structured feedback (rating scales, open-text) and visualizes aggregates.
The Discord bot also gives students instant feedback. Figure 2 shows how their responses are spread across satisfaction levels, making the overall sentiment transparent.
Figure 2. Immediate feedback results for students.
Survey Administration.
Instructors build complex or straightforward surveys, set time limits, and view the results analytics.
The bot offers two types of survey interactions in Discord channels: simple-question surveys and multi-step surveys. Figure 3 depicts a five-level difficulty survey suitable for rapid, in-lecture feedback, while Figure 4 presents the instructor’s real-time results.
Figure 3. Simple survey in a Discord channel.
Figure 4. Instructors’ results view.
For more detailed feedback, instructors can create complex surveys that begin in the channel and continue via direct messages after clicking the ”Participate” button, allowing a sequence of questions (see Figure 5). The results of these complex surveys are also presented to the instructors in similar charts, as shown in Figure 4.
Figure 5. Complex survey in a Discord channel.
4. Evaluation
The evaluation validated InsightEdu through a comparative study contrasting traditional CLI-based bot management with the mobile interface. Twenty experienced instructors with varying technical backgrounds performed six core tasks: attendance tracking, complex and straightforward survey deployment, feedback collection, configuration management, and historical data analysis.
Tests were conducted using a Raspberry Pi 4B over campus Wi-Fi to simulate realistic classroom environments. Usability was assessed using Jakob Nielsen’s heuristics, while cognitive load was measured via the NASA-TLX index. Quantitative metrics included task completion time, error rates, and system response latency. Qualitative assessment captured subjective satisfaction and interface preferences through structured interviews and rankings.
5. Performance and Usability Results
Quantitative analysis confirms that the InsightEdu mobile interface significantly outperforms traditional CLI workflows in efficiency, accuracy, and user satisfaction.
Efficiency and Error Rates. The mobile interface achieved a 40% reduction in median task completion time. The most substantial gains occurred in complex survey creation (65% faster) and data analysis (55% faster). By eliminating command-line syntax, error rates dropped from an average of 2.3 per task with CLI commands to almost none. While CLI errors were primarily syntax-related, in the app, they were essentially network connectivity issues.
Technical Performance. The system met all responsiveness targets over the campus Network. Latency metrics are summarized in Table 1. Database operations averaged 180 ms, and Discord API interactions for command execution averaged 250 ms. Analytics visualizations were generated with sub-100 ms latency.
Table 1. System Response Latency Metrics
Operation Type Average Latency
Local Chart Rendering << 100 ms
Database Operations 180 ms
Discord API Interaction 250 ms
Total Round-trip (95th percentile) << 300 ms
Usability and Cognitive Load. Heuristic evaluation yielded an average score of 4.2/5.0, with Aesthetic Design scoring highest (4.8/5.0). Qualitative feedback showed a 92% preference for the mobile application. NASA-TLX assessment indicated a significant reduction in perceived mental demand, as the interface moved from recall-based commands to recognition-based touch interactions.
Deployment Scalability. Stress tests on a Raspberry Pi 4B (4GB RAM) validated the system’s low-cost feasibility:
- Capacity: Supported 12 concurrent instructors managing 20–50 students each.
- Resources: Memory usage stabilized at 1.2GB; CPU utilization averaged 15–20%.
- Storage/Network: Bandwidth remained minimal (<<1KB per API request), and storage requirements were low ($\approx$50MB per semester).
- Reliability: Automatic reconnection restored full functionality within 30–60 seconds after network interruptions.
6. Discussion
This section analyzes the trade-offs between mobile-first design and the constraints of Discord-mediated instruction.
6.1. Implications for Educational Technology
The success of InsightEdu demonstrates that sophisticated learning analytics do not require expensive infrastructure. By combining free Discord hosting with low-cost Raspberry Pi hardware, the barrier to entry for interactive teaching is reduced by orders of magnitude compared to traditional LMS platforms. The 40% reduction in task completion time validates the importance of removing command-line friction for adoption among non-technical faculty. InsightEdu allows instructors to pivot from ”managing tools” to ”responsive pedagogy,” using real-time data to adjust lecture pacing.
6.2. Architectural Lessons and Design Patterns
While the MVVM pattern and dual-threaded Flask architecture provided high development speed, we identified some tensions:
Simplicity vs. Scalability: The single-process model is easy to deploy but introduces a single point of failure. Future iterations might require containerization for better fault isolation.
Security Pragmatism: Using static API keys (Level 2 REST) was appropriate for this cohort, though institutional scaling will necessitate transitioning to JWT or SSO integration.
Platform Synergy: Leveraging Discord’s reliable WebSocket infrastructure allowed us to focus on UX rather than protocol stability, confirming that ”building on proven platforms” is superior to custom-built chat solutions for EdTech.
6.3. Limitations and Threats to Validity
The study’s primary limitations include a small participant pool ($n=20$) and a lack of longitudinal data to track long-term adaptation. Furthermore, the system remains dependent on Discord’s API and Terms of Service. While manageable for pilots, institutional adoption requires contingency planning for platform-wide changes or service outages.
6.4. Future Research Directions
The logical evolution of InsightEdu involves three pillars:
AI Recommendations: Using LLMs to transform student feedback into actionable ”teaching tips” (e.g., ”70% of students found the last example difficult; suggest a recap”).
Cross-Platform Support: Moving toward Flutter or React Native to support Android users.
Pedagogical Outcomes: Shifting focus from ”instructor efficiency” to measuring actual student learning gains and engagement levels over a full semester.
7. Conclusion
InsightEdu bridges a critical usability gap by providing a touch-centric mobile interface for managing Discord-based educational bots. By decoupling bot logic from presentation through a native SwiftUI client and Flask API, the system eliminates the friction of command-line workflows. Evaluation results demonstrate a 40% reduction in task completion time, lower error rates, and a 92% instructor preference for the mobile interface over traditional CLI methods. Beyond usability, the successful deployment on low-cost Raspberry Pi hardware confirms that sophisticated learning analytics are viable in resource-constrained environments. By leveraging Discord’s free infrastructure, InsightEdu offers an accessible alternative to expensive, traditional LMS deployments. This work provides transferable design guidance for the next generation of EdTech: prioritizing mobile-first accessibility and operational simplicity is essential for achieving broad adoption and improving real-time pedagogical responsiveness in diverse academic settings.